A simple pendulum has a time period T1 on the earth surface and T2 when taken to height R abov e the Earth surface where R is the radius of the earth the value of T2/T1 is

Grade:

1 Answers

Arun
5 years ago
 
T=2pi*sqrt(l/g)
g is inversly proportional to (R+h)2 …..........h is height above the earth’s  surface
therefore T is proportional to R+h
T2/T1= (R+h)/R =2............ putting h=R